# PePNet: A Periodicity-Perceived Workload Prediction Network Supporting Rare Occurrence of Heavy Workload

### v2

@@ Abstract
Cloud providers can greatly benefit from accurate workload prediction. However, the workload of cloud servers is highly variable, with occasional heavy workload bursts. This makes workload prediction challenging. There are mainly two categories of workload prediction methods: statistical methods and neural-network-based ones. The former ones rely on strong mathematical assumptions and have reported low accuracy when predicting highly variable workload. The latter ones offer higher overall accuracy, yet they are vulnerable to data imbalance between heavy workload and common one. This impairs the prediction accuracy of neural network-based models on heavy workload. Either the overall inaccuracy of statistic methods or the heavy-workload inaccuracy of neural-network-based models can cause service level agreement violations. Thus, we propose PePNet to improve overall especially heavy workload prediction accuracy. It has two distinctive characteristics: (i) A Periodicity-Perceived Mechanism to detect the existence of periodicity and the length of one period automatically, without any priori knowledge. Furthermore, it fuses periodic information adaptively, which is suitable for periodic, lax periodic and aperiodic time series. (ii) An Achilles’ Heel Loss Function iteratively optimizing the most under-fitting part in predicting sequence for each step, which evidently improves the prediction accuracy of heavy load. Extensive experiments conducted on Alibaba2018, SMD dataset and Dinda’s dataset demonstrate that PePNet improves MSE for overall workload by 11.8% on average, compared with state-of-the-art methods. Especially, PePNet improves MSE for heavy workload by 21.0% on average.
Index Terms:
Time series forecasting, cloud computation.
@@ I Introduction
Accurate workload prediction brings huge economic benefits to cloud providers [1] and many cloud frameworks make real-time adjustments based on the workload prediction results [2]. On the one hand, workload prediction provides meaningful insights to improve the utilization of cloud servers while ensuring the quality of service (QoS) [3] [4]. On the other hand, it can also alarm the forthcoming uncommon heavy workload, thus helping to avoid service level agreement (SLA) violations.
@@
(c) The first half of Achilles’ Heel Loss Function fitting process.
(d) The second half of Achilles’ Heel Loss Function fitting process.
Figure 1: (a)-(b) Overall workload and heavy workload prediction accuracy. (c)-(d) The fitting process of Achilles’ Heel Loss Function. (c) In the first half of the process, the most under-fitting part is on the right. (d) After some training iterations, the prediction error on the right diminishes and the most under-fitting part shifts to the left. Then, Achilles’ Heel Loss Function pays more attention to the left part in this phase.
However, the high variability of cloud-server workload [5] [6] makes workload prediction challenging. Statistical methods and neural-network-based methods are two mainstream workload prediction methods. The former ones require the time series to satisfy strong mathematical assumptions and are not robust when predicting highly variable workload [5]. For example, the classical statistic method, ARIMA [7], requires the time series to be stationary after difference and shows dissatisfied results when predicting highly variable workload [5]. The neural-network-based methods are more suitable to predict highly variable workloads, but they are vulnerable to data imbalance between heavy workload and common one (i.e. the heavy workload is much more uncommon [8]). To demonstrate this, we present statistics of the heavy workload proportion of different datasets in Tab. I, where the heavy workload is defined as the workload greater than the average workload plus one standard deviation for each machine. It has also been proven that the data imbalance between heavy workload and common workload impairs the former’s accuracy [9]. As an empirical example, Fig. 1(a)-Fig. 1(b) shows the overall prediction accuracy and heavy-workload prediction accuracy of some prevailing methods (ARIMA [7], LSTM with attention (LSTMa) [10], Informer [11], Learning based Prediction Algorithm for cloud workload (L-PAW) [5], LSTM and GRU) on Alibaba2018 dataset, where the heavy-workload predicted error is nearly twice as large as the overall predicted error. The inaccurate prediction will not only reduce the utilization of cloud servers but also bring SLA violations, as it provides wrong information to the scheduler.
TABLE I: Proportion of heavy workload.
Dataset Heavy workload proportion
CPU usage of Alibaba’s dataset [^1] https://github.com/alibaba/clusterdata 15.73%
Memory usage of Alibaba’s dataset 12.89%
Workload of Dinda’s dataset [12] 11.15%
To predict overall workload and heavy workload accurately, we propose a Periodic-Perceived Workload Prediction Network (PePNet), which supports the prediction of highly variable workload with the rarely-occurring heavy workload. PePNet mainly improves the prediction accuracy in two aspects: 1) using the periodically recurrent patterns to guide the workload prediction and 2) using an Achilles’ Heel Loss Function [^2] Achilles’ Heel originates from Greek mythology and is the fatal weakness of the hero Achilles, which pays more attention to the under-fitting part to offset the negative influence of data imbalance.
The existing methods that embed the periodic information [13, 14, 15, 16] are applied to the situations where the periodicity is stable and known (i.e. whether the time series is periodic and the length of a period are fixed and known). However, the cloud environment is complex and highly dynamic. Different servers exhibit different periodicity and even the same server exhibits different periodicity at different time. This calls for a technique that detects periodicity and period length automatically and is capable of manipulating time series with mixed periodicity (i.e. the time series is periodic for some time and aperiodic for the other time). Thus, we propose a Periodicity-Perceived Mechanism, which consists of Periodicity-Mining Module and Periodicity-Fusing Module. The Periodicity-Mining Module detects the periodicity and period length and the Periodicity-Fusing Module adaptively fuses the periodic information for periodic and aperiodic time series. We mathematically prove the error bound of the periodic information extracted by Periodicity-Mining Module. Furthermore, we design an automatic hyperparameter determination method to provide convenience for the training process.
The Achilles’ Heel Loss Function specifically targets the upper bound of the prediction error, which enables PePNet to strengthen and improve the most vulnerable part of the prediction. At each training step, the Achilles’ Heel Loss Function picks the most under-fitting part in a predicting sequence and minimizes its prediction error. As shown in Fig.1(c)-1(d), when a kind of pattern is well learned, the most under-fitting part will shift to the position of another pattern. Thus, each part in the predicting sequence is fitted alternatively. This characteristic helps alleviate the problem of overemphasizing sporadic patterns, a limitation that has been observed in many loss functions designed to combat under-fitting issues [17]. In this way, Achilles’ Heel Loss Function solves the under-fitting problem of heavy load and improves the heavy-load prediction accuracy as well as the overall prediction accuracy.
The main contributions of our work are summarized as:
@@
  • We design a Periodicity-Perceived Mechanism, which mines and adaptively fuses the periodic information. Moreover, we mathematically prove an error bound of periodic information extracted by Periodicity-Mining Module and propose an automatic hyperparameter determining method for it.
  • We have formulated an Achilles’ Heel loss function aimed at enhancing the accuracy of heavy workload prediction and overall prediction performance. This approach effectively mitigates the issue of overemphasizing sporadic patterns, a common limitation faced by many loss functions designed to combat under-fitting problems.
@@
  • We conduct extensive experiments on the Alibaba2018, SMD and Dinda’s datasets and demonstrate that PePNet improves the heavy-workload prediction accuracy (MSE) by 21.0% on average, while promoting overall prediction accuracy (MSE) by 11.8% on average.
@@ II Methodology
Preliminary. The input of data $X$ is divided into three parts, which are denoted by $X^{enc}_{short}\in\mathbb{R}^{I\times d}$, $X^{enc}_{long}\in\mathbb{R}^{M\times d}$, $X_{period}\in\mathbb{R}^{P\times d}$ respectively. The $d$ is the dimension of the feature at each time slot and $I,M$ and $P$ respectively stand for the length of short-term dependent information, long-term tendency information and periodic information. The data division is shown in Fig. 3(a): $Y$ is the ground truth value of predicting workload; $X^{enc}_{short}$ is the nearest workload time series before the predicting part and shows high relevance to predicting workload; $X^{enc}_{long}$ is a bit longer workload time series before $X^{enc}_{short}$ and reflects the long term tendency in workload variation; $X_{period}$ is the first period of workload for each machine; The $y_{period}$ is the workload sequence corresponding to $Y$ in the period of $X_{period}$. The extracting process of $X_{period}$ and $y_{period}$ is illustrated in section II-B1. We use $X^{enc}_{long,i}$ and $X^{enc}_{short,i}$ to denote the $i$-th time slot’s workload in $X^{enc}_{long}$ and $X^{enc}_{short}$ respectively.
The three kinds of information (i.e. short-term dependent information, long-term tendency information and periodic information) bring different effects on workload prediction. The short-term dependent information $X^{enc}_{short}$ is highly related to the predicting sequence, as the customers’ behaviors are continuous and the workload is similar in a short time slot. The long-term tendency information denotes the forthcoming heavy load, as the heavy load is more likely to happen in ascending tendency. This tendency information is usually masked by noises in the short term and has to be extracted from long-term time series. As shown in Fig. 3(b), we pick a small segment of the workload from a long-term-ascending sequence. Though the long-term ascending tendency it is, the short-term workload just fluctuates around a stable value and does not show any increasing tendency. The periodic information provides the recurrent patterns in history and promotes both overall prediction accuracy and heavy-load prediction accuracy.
(a) The input data division.
(b) The long-term tendency information
(c) Local periodicity violation in lax periodic series.
(d) Autocorrelation coefficient for periodic sequence.
(e) Autocorrelation coefficient for aperiodic sequence.
(f) The method of computing periodic information.
(g) Histogram of values for the first peak of the autocorrelation coefficient.
(h) How the value of $\gamma$ impacts the distribution of gradient weights.
@@
II-A Overview of PePNet
PePNet is based on an encoder-decoder [18] architecture and fuses three kinds of information from rough data: the long-term tendency information, the short-term dependent information and the periodic information. Besides, the whole model is trained by an Achilles’ Heel Loss Function to offset the negative impact of data imbalance. The overview of PePNet is shown in Fig. 2.
In the encoder, PePNet uses a backbone encoder, which can be implemented as many popular time-series-manipulating structures, such as transformer [19], LSTM [20] etc., to capture short-term dependent information from $X^{enc}_{short}$ and uses Periodicity-Mining Module to detect and extract periodic information from $X_{period}$. As for long-term tendency information, PePNet firstly downsamples $X^{enc}_{long}$, as it helps to improve the processing efficiency as well as maintain and magnify the trend information of time series. Then, PePNet uses self-attention to manipulate the down-sampled sequence. The processes of extracting short-term, long-term and periodic information are shown in Eq.1-3. In function $\operatorname{Attention}(q,k,v)$, $q$, $k$, $v$ respectively stand for the query, key and value[21].
$\overline{X}^{enc}_{short,i+1}=\operatorname{Backbone-Encoder}(X^{enc}_{short,i})$ (1)
$y_{period},X_{period}=\operatorname{PeriodicityMining}(X,X^{enc}_{short})$ (2)
$\overline{X}^{enc}_{long}=\operatorname{Downsample}(\operatorname{Attention}(X^{enc}_{long},X^{enc}_{long},X^{enc}_{long}))$ (3)
In the decoder, PePNet uses a backbone-decoder, which can be implemented as many popular time-series-manipulating structures, such as transformer, LSTM, etc., to generate $\hat{y}\in\mathbb{R}^{J\times d}$, where $J$ stands for the predicting length. When using recurrent neural networks such as LSTM or GRU, the concatenation of $h_{j}$ in Eq. 4 serves as the hidden state, and $\hat{y}_{j}$ serves as the network input. When using a transformer or its variants, $\hat{y}_{j}$ serves as the query and $h_{j}$ serves as both the keys and values. After that, PePNet uses Periodicity-Fusing Module to estimate the reliability of $\hat{y}$ and fuses it with $y_{period}$ according to their reliability. By then, the Periodicity-Fusing Module generates the final prediction $y$. The process of decoder is shown in Eq.4-6, where $W_{h},W_{c}$ are both model parameters.
$\hat{y}_{j+1}=\operatorname{Backbone-Decoder}(\hat{y}_{j},h)$ (4)
$h=\operatorname{Attention}(X^{enc}_{short},\overline{X}^{enc}_{long},\overline{X}^{enc}_{long})$ (5)
$y=\operatorname{PeriodicityFusing}(X^{enc}_{short},y_{period},\hat{y})$ (6)
When training the model, the imbalance between heavy load and normal workload usually leads to unsatisfactory accuracy of heavy load prediction. Thus, we propose an Achilles’ Heel Loss function, which aims to minimize the upper bound of prediction error. At each step, Achilles’ Heel Loss Function iteratively optimizes the most under-fitting part for this step, in order to improve the prediction accuracy of the heavy load when keeping the overall prediction accurate. It is named after Achilles’ Heel as it focuses on the most vulnerable portion along the predicting sequence. Thus, the prediction accuracy of the heavy load becomes more accurate, when keeping the overall prediction accurate.
II-B Periodicity-Perceived Mechanism
Periodic information can effectively improve the overall prediction accuracy as well as the heavy workload prediction accuracy. It consists of Periodicity-Mining Module and Periodicity-Fusing Module.
There are three main challenges to fuse periodic information. (1) We have no priori knowledge about the periodicity of workload for different machines, which calls for the Periodicity-Mining Module to detect the existence of periodicity and the length of one period. (2) The periodicity for different machines is variable (i.e., strict periodic, lax periodic, aperiodic), which calls for an adaptive Periodicity-Fusing Module. (3) It is hard to fuse the periodic information in lax periodic series because lax periodic series are sometimes periodic and sometimes not. In Fig. 3(c), we overlap a series shifted one period ahead and the original series. As Fig. 3(c) shows, there are mainly two obstacles to fuse periodic information in lax periodic series: periodic shift and local periodicity violation. The former one can be solved by dynamic matching which is depicted in section Periodicity-Mining Module. The latter is caused by noise and external events. Therefore, we filter out the noise in periodic information in Periodicity-Mining Module and evaluate the reliability of periodic information in the Periodicity-Fusing Module.
II-B1 Periodicity-Mining Module
PePNet calculates the time series autocorrelation coefficient $\rho_{k}$ for each machine as shown in Eq.7, which represents the linear correlation among all workloads with interval $k$ (i.e., linear correlation between $X_{t}$ and $X_{t-k}$, for every $t\in\{t\in\mathbb{N}^{+}|t<sequence\ length\}$, $X_{t}$ stands for the workload at $t$-th time slot). As shown in Fig. 3(d), when the workload is periodic, the autocorrelation coefficient rises to a large value again after the first decline, which represents a high hop relevance of workload. While the sequence is aperiodic, the autocorrelation coefficients are shown in Fig. 3(e), which have a distinct pattern. Therefore, PePNet sets a hyperparameter $\mathcal{T}$ and judges whether a time series is periodic by detecting whether the autocorrelation coefficient crosses over $\mathcal{T}$ again. The value of $\mathcal{T}$ denotes the acceptable range of periodicity strictness.
$\rho_{k}=\frac{cov(X_{t-k},X_{t})}{\sqrt{cov(X_{t},X_{t})cov(X_{t-k},X_{t-k})}}$ (7)
The position of the first peak in autocorrelation coefficients of periodic sequence strikes the highest hop relevance and denotes the length of one period. We illustrate it in Fig. 3(d). The reason why the autocorrelation coefficient shows such a pattern is that the user behavior is continuous in time, so the autocorrelation coefficients of small $k$ are large. As the time interval $k$ increases, the relevance drops down. But when the time interval $k$ reaches around the integer multiple of period length, the workload at the two moments $X_{t}$ and $X_{t-k}$ shows a high linear correlation again.
Based on these observations, as shown in Algorithm 1, PePNet traverses the value of $k$ from small to large. If PePNet finds the smallest $k$ that makes the autocorrelation coefficient $\rho_{k}$ satisfy the conditions: $\rho_{k}>\rho_{k-1},\rho_{k}>\rho_{k+1}$ and $\rho_{k}>\mathcal{T}$, the time series is periodic and the length of one period is $k$. Otherwise, the time series is aperiodic. We cut out the first period of the machines whose workloads are periodic in the training set as a periodic information knowledge base, which is denoted by $X_{period}$. To solve the issue of period shift, PePNet uses a dynamic matching. If the workload is periodic, PePNet finds the sequence $\{X_{t_{a}},\ldots,X_{t_{a}+I-1}\}$ in $X_{period}$, whose distance from $X^{enc}_{short}$ is the smallest. The distance can be computed either by Mean Square Error (MSE) or by Dynamic Time Warping (DTW). Then, PePNet sets $y_{period}=\{X_{t_{a}+I},\ldots,X_{t_{a}+I+J}\}$. Otherwise, $\{-1,-1,\ldots,-1\}$ is set to $y_{period}$. Such a process is illustrated in Fig. 3(f).
Data: Workload time series $X$, Short-term dependent information $X^{enc}_{short}$
Result: The periodic information for prediction $y_{period}$, The periodic knowledge base $X_{period}$
1 $k\leftarrow-1$;
2 while $k<length(X)$ do
3      $k\leftarrow k+1$;
4       Compute $\rho_{k}$;
5       if $\rho_{k}>\rho_{k-1}\ and\ \rho_{k}>\rho_{k+1}\ and\ \rho_{k}>\mathcal{T}$ then
6            Period length=k ;
7             FoundPeriodLength=True ;
8             break ;
9       end if
10
11 end while
12if FoundPeriodLength then
13       $X_{period}=X[:Period\ length]$;
14       $y_{pos}=\arg\min_{t_{a}}Distance(X[t_{a}:t_{a}+I],X^{enc}_{short})$;
15       $y_{period}=X[y_{pos}:y_{pos}+I]$;
16
17 end if
18else
19       $X_{period}=-1$;
20       $y_{period}=[-1,\ldots,-1]$;
21
22 end if
Algorithm 1 Periodicity-Mining Module
@@ 20       $y_{period}=[-1,\ldots,-1]$;
Error estimation. We use $E[(X_{t-k}-X_{t})^{2}]$ to measure the quality of periodic length $k$ found by the Periodicity-Mining Module. We have the following result, where $\sigma_{t-k},\sigma_{t},\mu_{t-k},\mu_{t}$ are the standard deviation and expectation at $t-k$ and $t$ time slots respectively. The $\Delta\sigma$ and $\Delta\mu$ are $\sigma_{t-k}-\sigma_{t}$ and $\mu_{t-k}-\mu_{t}$ respectively. $\sigma$ is the standard deviation of a stationary time series.
@@
Theorem 1. When using Algorithm 1, $E[(X_{t-k}-X_{t})^{2}]$ is upper bounded by $(\Delta\sigma)^{2}+(\Delta\mu)^{2}+2(1-\mathcal{T})\sigma_{t-k}\sigma_{t}$. Furthermore, when the time series is stationary, $E[(X_{t-k}-X_{t})^{2}]$ is upper bounded by $2(1-\mathcal{T})\sigma^{2}$.
Proof of Theorem 1. According to algorithm 1, when PePNet finds $k$, Eq.8-9 holds, where $a_{i}=X_{i}-\mu_{i}$, $\tilde{t}=t-k$. $E[(X_{\tilde{t}}-X_{t})^{2}]$ can be transformed to Eq.10. Then, we substitute Eq.9 into Eq.10 and obtain Eq.11. The Eq.11 can be further reduced into Eq.12 and the first part of Theorem 1 is proven. When the time series is stationary, $\Delta\sigma^{2}$ and $\Delta\mu^{2}$ are zero. Then the upper bound can be reduced to $2(1-\mathcal{T})\sigma_{t-k}\sigma_{t}$.
$\frac{E[a_{\tilde{t}}a_{t}]}{\sigma_{\tilde{t}}\sigma_{t}}>\mathcal{T}$ (8)
$E[a_{\tilde{t}}a_{t}]>\mathcal{T}*\sigma_{\tilde{t}}\sigma_{t}$ (9)
@@
Taking a further look at the upper bound, the first and second items of the upper bound are constants. Thus, the value of error depends on the last item, which is determined by the value of $\mathcal{T}$. Considering that the $\mathcal{T}$ is less than or equal to one, the closer the $\mathcal{T}$ is to one, the smaller the error is. However, most of the time series are lax periodic in reality. Setting the $\mathcal{T}$ as 1 may miss many lax periodic time series. Thus, there is a tradeoff between improving the accuracy of periodic length and finding all the periodic time series (including lax periodic time series).
@@
Considering the tradeoff may make it challenging to choose a proper value of $\mathcal{T}$, we provide an automatic setting method for $\mathcal{T}$ based on statistical observations. For finer tuning, this method provides a searching range for $\mathcal{P}$. The hyperparameter $\mathcal{P}$ is a threshold setting for the first peak of the autocorrelation coefficient. Thus, we look into the distributions of the first peak value. Taking Alibaba2018-CPU [^3] https://github.com/alibaba/clusterdata as an example, we take 75 machines in it and plot a histogram of first-peak value in Fig. 3(g). There are two properties of the distribution: clustered and following Gaussian Distribution for each cluster. We use the Gaussian Mixture Model (GMM) algorithm to fit the mixed Gaussian distribution of its first peak values and plot the Gaussian distributed curve on the histogram in Fig. 3(g). The parameter for this GMM is given in Tab. II, where ”std” is the shorthand for standard deviation. Different clusters have different degrees of periodicity. The higher the expectation of the cluster is, the stricter its periodicity is. To guarantee the quality of the periodicity the Periodicity-Mining Module detects, it picks the workload in the cluster with the highest expected value as periodic and filters the others. Let $\mu$ and $\sigma$ denote the expectation and standard deviation of the cluster with the highest expectation. In industrial applications, $\mu-\sigma$ can be directly used as the $\mathcal{P}$ to save the labor costs, as the probability that the first peak is greater than $\mu-\sigma$ in this cluster is 84.2% according to Gaussian distribution. For finer tuning, the $mathcal{P}$ can be explored between $\mu-\sigma$ and $\mu+\sigma$.
TABLE II: Parameters of Gaussian distribution.
weight expectation std
Gaussian distribution 1 0.84 0.53 0.06
Gaussian distribution 2 0.10 0.23 0.03
Gaussian distribution 3 0.06 -0.04 0.05

### v3

# Feiyi Chen1, Naijin Liu2, Zhen Qin1, Hailiang Zhao1, Mengchu Zhou3, Shuiguang Deng1

@@ Abstract
Cloud providers can greatly benefit from accurate workload prediction. However, the workload of cloud servers is highly variable, with occasional workload bursts, which makes workload prediction challenging. The time series forecasting methods relying on periodicity information, often assume fixed and known periodicity length, which does not align with the periodicity-changeable nature of cloud service workloads. Although many state-of-the-art time-series forecasting methods do not rely on periodicity information and achieve high overall accuracy, they are vulnerable to data imbalance between heavy workloads and regular workloads. As a result, their prediction accuracy on rare heavy workloads is limited. Unfortunately, heavyload-prediction accuracy is more important than overall one, as errors in heavyload prediction are more likely to cause Service Level Agreement violations than errors in normal-load prediction. Thus, we propose a changeable-periodicity-perceived workload prediction network (PePNet) to fuse periodic information adaptively for periodicity-changeable time series and improve rare heavy workload prediction accuracy. It has two distinctive characteristics: (i) A Periodicity-Perceived Mechanism to detect the periodicity length automatically and fuses periodic information adaptively, which is suitable for periodicity-changeable time series, and (ii) An Achilles’ Heel Loss Function that is used to iteratively optimize the most under-fitting part in predicting sequence for each step, thus evidently improving the prediction accuracy of heavy load. Extensive experiments conducted on real-world datasets demonstrate that PePNet improves accuracy for overall workload by 11.8% averagely, compared with state-of-the-art methods. Especially, PePNet improves accuracy for heavy workload by 21.0% averagely.
@@ I Introduction
AccurateAccurate workload prediction brings huge economic benefits to cloud providers [29] and many cloud frameworks make real-time adjustments based on the workload prediction results [2]. On the one hand, workload prediction provides meaningful insights to improve the utilization of cloud servers while ensuring the Quality of Service (QoS) [15] [28, 4, 3, 39]. On the other hand, it can alarm forthcoming rare heavy workload to avoid Service Level Agreement (SLA) violations.
However, the high variability of cloud-server workload [8, 1] makes workload prediction challenging. For time-series forecasting methods that rely on periodicity information [17, 6, 25, 37], they usually assume fixed and known periodic length, which is not in line with dynamic cloud environments where periodic patterns vary across services and over time. Those SOTA methods that do not highly rely on periodicity [34, 10, 21, 16, 44, 30, 23] show better generalization performance than those do so. However, they are vulnerable to data imbalance between heavy workload and common one and have been reported much lower performances on rare heavy workload burst [20, 13]. As an empirical example, Fig. 1(a)- 1(b) show the overall prediction accuracy and heavy-workload one of some prevailing methods (ARIMA [5], LSTM with attention (LSTMa) [45], Informer [42], Learning based Prediction Algorithm for cloud workload (L-PAW) [8], LSTM and GRU), where the latter’s error is nearly twice as large as the former’s error. However, accurately predicting heavy workloads is more important than predicting regular ones, because errors in heavy-workload prediction are more likely to cause SLA violations than those in normal-load prediction, as they provide misleading information to the scheduler during workload bursts.
To cope with time series with changeable periodicity and to accurately predict heavy workloads, we propose PePNet, which enables accurate prediction of highly variable workloads, including rarely occurring heavy-workload events. It improves the prediction accuracy in two aspects: 1) mining and fusing periodic information adaptively when periodicity is changeable without any priori knowledge; and 2) using an Achilles’ Heel Loss Function (AHLF)  [^1] Achilles’ Heel originates from Greek mythology and is the fatal weakness of the hero Achilles, which pays more attention to the under-fitting part to offset the negative influence of data imbalance. To address the problem of changeable periodicity in cloud environment, we propose a Periodicity-Perceived Mechanism with a Periodicity-Mining Module for automatic periodicity and period-length detection and a Periodicity-Fusing Module for adaptively handling mixed periodicity. We further provide a theoretical error bound for the mined periodic information and an automatic hyperparameter selection strategy. AHLF focuses on the upper bound of prediction error by iteratively minimizing the most under-fitted segment in a prediction sequence. This alternating fitting strategy mitigates the overemphasis on sporadic patterns observed in many loss functions combatting the under-fitting problem [14], thereby alleviating under-fitting under heavy load and improving heavy-load and overall prediction accuracy.
@@
Figure 1: (a)-(b) Overall workload and heavy workload prediction accuracy
We aim to make the following novel contributions to the field of workload prediction in cloud computing:
@@
  • We propose a Periodicity-Perceived Mechanism that mines and adaptively fuses periodic information for time series with changeable periodicity, with a theoretical error bound and an automatic hyperparameter selection strategy.
  • We formulate an AHLF to improve heavy-workload and overall prediction accuracy by mitigating the overemphasis on sporadic patterns.
@@
  • We perform extensive experiments on Alibaba2018, Server MAchine Dataset (SMD), and Dinda’s datasets to show that PePNet improves heavy-workload Mean Square Error (MSE) by 21.0% and overall MSE by 11.8% on average.
@@ II Methodology
Preliminary. Input data $X$ is divided into three parts, which are denoted by $\bar{X}\in\mathbb{R}^{I\times d}$, $\tilde{X}\in\mathbb{R}^{M\times d}$, $X_{p}\in\mathbb{R}^{P\times d}$ respectively. The $d$ is the dimension of a feature vector at each time slot and $I,M$ and $P$ respectively stand for the length of short-term dependent information, long-term tendency information and periodic information. $\bar{X}$ is the nearest workload time series before the predicting part and shows high relevance to predicting workload; $\tilde{X}$ is a bit longer workload time series before $\bar{X}$ and reflects the long term tendency in workload variation; $X_{p}$ is the first period of workload for each machine; $Y$ is the ground truth value of predicted workload. We use $\tilde{X}_{i}$ and $\bar{X}_{i}$ to denote the $i$-th time slot’s workload in $\tilde{X}$ and $\bar{X}$ respectively.
II-A Overview of PePNet
PePNet is based on an encoder-decoder [9] architecture, which can be implemented as a transformer, LSTM, etc. It fuses three kinds of information from rough data: long-term tendency, the short-term dependent and periodic information. Besides, the whole model is trained by minimizing AHLF to offset the negative impact of data imbalance. The overview of PePNet is shown in Fig. 2.
In the encoder, the processes of extracting short-term, long-term and periodic information are shown as follows. In function $\operatorname{Attention}(q,k,v)$, $q$, $k$, and $v$ respectively stand for the query, key and value[32]. PeriodicityMining module is introduced in subsection II-B. The downsample module downsamples the input time series using an increasing arithmetic sequence as the stride.
$\dot{\bar{X}}_{i+1}=\operatorname{Encoder}(\bar{X}_{i})$ (1)
$y_{p},X_{p}=\operatorname{PeriodicityMining}(X,\bar{X})$ (2)
$)\dot{\tilde{X}}=\operatorname{Downsample}(\operatorname{Attention}(\tilde{X},\tilde{X},\tilde{X}$)) (3)
The process of decoder is given as follows, where $W_{h},W_{c}$ are both model parameters and the PreiodicityFusing module is introduced in subsection II-B.
$\hat{y}_{j+1}=\operatorname{Decoder}(\hat{y}_{j},h)$ (4)
[…]
II-B Periodicity-Perceived Mechanism
Periodic information can be used to effectively improve the overall prediction accuracy as well as heavy workload. It consists of Periodicity-Mining Module and Periodicity-Fusing Module.
II-B1 Periodicity-Mining Module
PePNet calculates the time series autocorrelation coefficient $\rho_{k}$ for each machine as:
$\rho_{k}=\frac{\operatorname{cov}(X_{t-k},X_{t})}{\sqrt{\operatorname{cov}(X_{t},X_{t})\operatorname{cov}(X_{t-k},X_{t-k})}}$ (7)
which represents the linear correlation among all workloads with interval $k$ (i.e., linear correlation between $X_{t}$ and $X_{t-k}$, for every $t\in\{t\in\mathbb{N}^{+}|t<l\}$, $X_{t}$ stands for the workload at the $t$-th time slot). As shown in Fig. 3(a), when the workload is periodic, the autocorrelation coefficient rises to a large value again after the first decline, which represents a high hop relevance of workload. While the sequence is aperiodic, the autocorrelation coefficients are shown in Fig. 3(b), which have a distinct pattern. Therefore, PePNet sets a hyperparameter $\mathcal{T}$ and judges whether a time series is periodic by detecting whether the autocorrelation coefficient crosses over $\mathcal{T}$ again. The value of $\mathcal{T}$ denotes the acceptable range of periodicity strictness.
The position of the first peak in autocorrelation coefficients of periodic sequence strikes the highest hop relevance and denotes the length of one period. We illustrate it in Fig. 3(a). The reason why the autocorrelation coefficient shows such a pattern is that the user behavior is continuous in time, thus autocorrelation coefficients of small $k$ are large. As $k$ increases, the relevance drops down. But when $k$ reaches around the integer multiple of period length, the workload at the two moments $X_{t}$ and $X_{t-k}$ shows a high linear correlation again.
Based on these observations, PePNet traverses the value of $k$ from small to large. If PePNet finds the smallest $k$ that makes the autocorrelation coefficient $\rho_{k}$ satisfy the conditions: $\rho_{k}>\rho_{k-1},\rho_{k}>\rho_{k+1}$ and $\rho_{k}>\mathcal{T}$, the time series is periodic and the length of one period is $k$. Otherwise, it is aperiodic. We cut out the first period of the machines whose workloads are periodic in the training set as a periodic information knowledge base, which is denoted by $X_{p}$. To solve the issue of period shift, PePNet uses a dynamic matching. If the workload is periodic, PePNet finds the sequence $\{X_{t_{a}},\ldots,X_{t_{a}+I-1}\}$ in $X_{p}$, whose distance from $\bar{X}$ is the smallest. The distance can be computed either by MSE or by Dynamic Time Warping (DTW). Then, PePNet sets $y_{p}=\{X_{t_{a}+I},\ldots,X_{t_{a}+I+J}\}$. Otherwise, $\{-1,-1,\ldots,-1\}$ is set to $y_{p}$.
(a) Autocorrelation coefficient for periodic sequence.
(b) Autocorrelation coefficient for aperiodic sequence.
(c) Histogram of values for the first peak of the autocorrelation coefficient.
(d) How the value of $\gamma$ impacts the distribution of gradient weights.
@@ II-B1 Periodicity-Mining Module
Error estimation. We use $E[(X_{t-k}-X_{t})^{2}]$ to measure the quality of periodic length $k$ found by Periodicity-Mining Module. We have the following result, where $\sigma_{t-k},\sigma_{t},\mu_{t-k},$ and $\mu_{t}$ are the standard deviation and expectation at tkt-k and $t$ time slots respectively, $\Delta\sigma=\sigma_{t-k}-\sigma_{t}$ and $\Delta\mu=\mu_{t-k}-\mu_{t}$. $\sigma$ is the standard deviation of a stationary time series.
@@
Theorem 1. When using Periodicity-Mining Module, $E[(X_{t-k}-X_{t})^{2}]$ is upper-bounded by $(\Delta\sigma)^{2}+(\Delta\mu)^{2}+2(1-\mathcal{T})\sigma_{t-k}\sigma_{t}$. Furthermore, when the time series is stationary, $E[(X_{t-k}-X_{t})^{2}]$ is upper-bounded by $2(1-\mathcal{T})\sigma^{2}$.
Proof of Theorem 1 is given in Appendix A.
@@
Taking a further look at the upper bound, its first and second items are constants. Thus, the value of error depends on the last item, which is determined by the value of $\mathcal{T}$. Considering that $\mathcal{T}$ is less than or equal to one, the closer $\mathcal{T}$ is to one, the smaller the error is. However, most of the time series are lax-periodic in reality. Setting $\mathcal{T}$ as 1 may miss many lax-periodic time series. Thus, there is a tradeoff between improving the accuracy of periodic length and finding all the periodic time series (including lax-periodic ones).
@@
Considering that tradeoff may make it challenging to choose a proper value of $\mathcal{T}$, we provide an automatic setting method for $\mathcal{T}$ based on statistical observations. For finer tuning, this method provides a searching range for autocorrelation coefficient first peak threshold $\mathcal{H}$. Taking Alibaba2018-CPU [^2] https://github.com/alibaba/clusterdata as an example, we take 75 machines in it and plot a histogram of first-peak value in Fig. 3(c). There are two properties of the distribution: clustered and following Gaussian Distribution for each cluster. We use Gaussian Mixture Model (GMM) [33] to fit the mixed Gaussian distribution of its first peak values and plot the Gaussian distributed curve on the histogram in Fig. 3(c). Different clusters have different degrees of periodicity. The higher the expectation of the cluster is, the stricter its periodicity is. To guarantee the quality of the periodicity that Periodicity-Mining Module detects, it picks the workload in the cluster with the highest expected value as periodic and filters the others. Let $\mu$ and $\sigma$ denote the expectation and standard deviation of the cluster with the highest expectation. In industrial applications, $\mu-\sigma$ can be directly used as $\mathcal{H}$ to save the labor cost, as the probability that the first peak is greater than $\mu-\sigma$ in this cluster is 84.2% according to Gaussian distribution. For finer tuning, $\mathcal{H}$ can be explored between $\mu-\sigma$ and $\mu+\sigma$.